# Convergence and error control of consistent PINNs for elliptic PDEs

### v1

@@ 4.1 Using FEMs or AFEMs
Let $\widehat{u}_{n}$ be the approximation to $\widetilde{u}$ obtained from the Galerkin projection onto a standard finite element space $V_{n}$ of dimension $n$. That is, $\widehat{u}_{n}$ is the Galerkin solution to (1) with right side $\widehat{f}$ and boundary value $\widehat{g}$. Then, we have
$\|u-\widehat{u}_{n}\|_{H^{1}(\Omega)}\lesssim R_{m}^{*}({\cal U})_{H^{1}(\Omega)}+\varepsilon_{n}({\cal U})_{H^{1}(\Omega)},$ (4.2)
where
$\varepsilon_{n}({\cal U})_{H^{1}(\Omega)}:=\mathop{\rm dist}({\cal U},V_{n})_{H^{1}(\Omega)},\quad n\geq 1.$ (4.3)
The term $\varepsilon_{n}$ will tend to zero as $n\to\infty$ and we would obtain the optimal recovery error on the right side of (4.2) if $n$ is chosen suitably large. The actual decay rate of $\varepsilon_{n}$ depends on the regularity of the model class ${\cal U}$ in $H^{1}(\Omega)$. Therefore, the issue becomes what do our model class assumptions $f\in{\cal F}$, $g\in{\cal G}$, say about the regularity of $u$ in the $H^{1}(\Omega)$ norm. If we use linear spaces such as standard finite element spaces, then we would want to know the regularity of $u$ in the scale of $H^{t}(\Omega)$ spaces, $t>1$. There are several theorems that obtain such regularity results (see e.g. [11, 35, 46] and the references therein) and thereby obtain concrete bounds on the decay rate $\varepsilon_{n}$. We do not elaborate on this further but refer the interested reader to the papers cited above.
If we use nonlinear numerical methods such as AFEMs from for solving (1) with right side $\widehat{f}$, and boundary value $\widehat{g}$, then the decay rate for $\varepsilon_{n}({\cal U})$ can be improved (see [5] for a summary of such results).
@@ 4.2 Reduced models
If one is faced with solving several PDEs with the same data sites $({\cal X},{\cal Z})$ but with different data $({\bf f},{\bf g})$, one can build a numerical algorithm as follows. Suppose that ${\cal X}=G_{k,r}$, $|G_{k,r}|=\widetilde{m}=[r2^{k}]^{d}$, see (3.8) and ${\cal Z}=\overline{G}_{k,r}$, $|\overline{G}_{k,r}|=\overline{m}=2d[r2^{k}]^{d-1}$, see (3.16) $\overline{\Omega}$ with spacing $h=r^{-1}2^{-k}$.
We know that the simplicial interpolation operator $S_{k}^{*}$ defined in (2.16) provides a near optimal recovery for ${\cal F}$. Namely,
@@
Similarly, the interpolant $\overline{S}_{k}(g)$ to $g\in{\cal G}$ used for optimal recovery, see (3.18), can be written as
$\overline{S}_{k}(g)=\sum_{i=1}^{\overline{m}}g({\bf z}_{i})\psi_{i},\quad\overline{m}=2d[r2^{k}]^{d-1}.$ (4.6)
@@
is a near optimal recovery of the solution $u$ to (1) with right side $f$ and boundary values $g$. Note that $\widetilde{u}$ is the solution to (1) with right side $S_{k}^{*}(f)$ and boundary value $\overline{S}_{k}(g)$.
@@
$\widehat{u}:=\sum_{i=1}^{\widetilde{m}}f({\bf x}_{i})\widehat{\phi}_{i}+\sum_{i=1}^{\overline{m}}g({\bf z}_{i})\widehat{\psi}_{i}$ (4.8)
@@ 5 Numerical methods based on optimization
Another approach to finding a near optimal approximation to $u$ from the given data is through optimization, such as the one used in PINNS. To analyze this approach, we begin this section by first considering a theoretical optimization algorithm which we shall later turn into a numerical procedure. We begin by introducing the functional ${\cal L}_{T}:H^{1}(\Omega)\to\mathbb{R}$, defined as
${\cal L}_{T}(v):=\|f+\Delta v\|_{H^{-1}(\Omega)}+\|v-g\|_{H^{1/2}(\partial\Omega)},$ (5.1)
for fixed $f\in{\cal F}\subset H^{-1}(\Omega)$ and $g\in{\cal G}\subset H^{1/2}(\partial\Omega)$. Clearly, the unique solution $u\in H^{1}(\Omega)$ to (1) with right side $f$ and boundary condition $g$ is the unique minimizer of this functional since ${\cal L}_{T}(u)=0$. In other words, we have
$u=\mathop{\rm argmin}_{v\in H^{1}(\Omega)}{\cal L}_{T}(v).$ (5.2)
Note also that because of (1.3), we have for every $v\in H^{1}(\Omega)$
$\|u-v\|_{H^{1}(\Omega)}\asymp\|f+\Delta v\|_{H^{-1}(\Omega)}+\|g-v\|_{H^{1/2}(\partial\Omega)}={\cal L}_{T}(v),$ (5.3)
with absolute constants in the equivalence. In other words, the loss ${\cal L}_{T}(v)$ for any $v$ tells us how close $v$ is to the true solution $u$.
The minimization in (5.2) is over all $v\in H^{1}(\Omega)$. To make this minimization more amenable to numerical implementation, we take this minimization over a set $\Sigma=\Sigma_{n}$, where $\Sigma$ is either a linear space of dimension $n$ or a nonlinear manifold of order $n$ (i.e., depending on $n$ parameters), and study any function
@@
$\|u-S_{\Sigma}\|_{H^{1}(\Omega)}\lesssim{\cal L}_{T}(S_{\Sigma})\lesssim{\cal L}_{T}(S)\lesssim\|u-S\|_{H^{1}(\Omega)},$ (5.5)
where the last inequality is a consequence of the left inequality in (1.3). Since (5.5) holds for all $S\in\Sigma$, we have that
@@ 6 Discretizing norms
The optimization procedure of the previous section is not a numerical algorithm since it does not incorporate numerical methods for estimating the norms appearing in the loss ${\cal L}_{T}$, namely $\|\cdot\|_{H^{-1}(\Omega)}$ and $\|\cdot\|_{H^{1/2}(\partial\Omega)}$. In this section, we address this issue by introducing discrete norms which involve only the values of functions at the data sites. Replacing the norms in ${\cal L}$ by these discrete norms leads to a loss which can be numerically computed. The equivalence between the norms in ${\cal L}_{T}$ and their discrete counterparts holds (modulo corresponding optimal recovery rates) for functions whose Laplacians are uniformly bounded in $\cal B$ and whose traces are uniformly bounded in $\overline{\cal B}$.
@@ 6.1 A discrete $L_{\tau}$ norm
In this section, we study the discretization of the $L_{\tau}$ norm for any $1\leq\tau\leq\infty$. Let ${\cal F}=U(B_{q}^{s}(L_{p}(\Omega)))$, $s>d/p,$ be a model class assumption on $f$. We consider the uniform grid $G_{k,r}\subset\overline{\Omega}$, see (3.8), consisting of $m$ points, $m=|G_{k,r}|=[r2^{k}]^{d}$, $k\geq 0$ and $r>\max\{s,1\}$. For any continuous function $f$ and any $1\leq\tau<\infty$, we define
$\|f\|^{*}_{L_{\tau}}:=\left[\frac{1}{m}\sum_{j=1}^{m}|f({\rm\bf x}_{j})|^{\tau}\right]^{1/\tau},\quad\hbox{where}\quad{\bf x}_{j}\in G_{k,r}.$ (6.1)
@@
$\|f\|_{L_{\tau}(\Omega)}\lesssim\|f\|^{*}_{L_{\tau}}+\|f\|_{{\cal B}}m^{-\alpha_{\tau}},\quad\hbox{and}\quad\|f\|^{*}_{L_{\tau}}\lesssim\|f\|_{L_{\tau}(\Omega)}+\|f\|_{{\cal B}}m^{-\alpha_{\tau}},$
with $\alpha_{\tau}:=\frac{s}{d}-\left(\frac{1}{p}-\frac{1}{\tau}\right)_{+}$. Recall that $m^{-\alpha_{\tau}}$ is the uniform rate of optimal recovery of the model class ${\cal F}$ in $L_{\tau}(\Omega)$. The constant in $\lesssim$ depends only on $r$ and $d$.
@@
We prove the lemma for $1\leq\tau<\infty$. A simple modification of the proof handles the case $\tau=\infty$. We use the simplicial interpolation operator $S^{*}_{k}$ given in (2.16). Let ${\cal T}_{k}:={\cal T}_{k}(\Omega)$ be the set of simplices that make up the simplicial decomposition of the cubes in ${\cal D}_{k}(\overline{\Omega})$. Then, from (2.16) we have
@@
where $L_{T}(f)$ is the polynomial in ${\cal P}^{d}_{r}$ that interpolates the data $\{f({\bf x}_{i}):\,{\bf x}_{i}\in\overline{T}\}$. By equivalence of norms on ${\cal P}^{d}_{r}$, we know that
$\|L_{T}(f)\|_{L_{\tau}(T)}\asymp\Bigl{[}\frac{1}{m}\sum_{{\bf x}_{i}\in{\overline{T}}}|f({\bf x}_{i})|^{\tau}\Bigr{]}^{1/\tau},$ (6.3)
@@
$\|S_{k}^{*}(f)\|_{L_{\tau}(\Omega)}\asymp\Bigl{[}\frac{1}{m}\sum_{i=1}^{m}|f({\bf x}_{i})|^{\tau}\Bigr{]}^{1/\tau}=\|f\|^{*}_{L_{\tau}(\Omega)}.$ (6.4)
For $f\in{\cal B}=B_{q}^{s}(L_{p}(\Omega)$, the piecewise polynomial interpolant $S_{k}^{*}(f)$ provides a near optimal recovery of $f$ from the data ${\bf f}$, that is
$\|f-S_{k}^{*}(f)\|_{L_{\tau}(\Omega)}\leq C\|f\|_{{\cal B}}m^{-\alpha_{\tau}}.$ (6.5)
It follows then from (6.4) that
$\left|\|f\|_{L_{\tau}(\Omega)}-\|S_{k}^{*}(f)\|_{L_{\tau}(\Omega)}\right|\leq C\|f\|_{{\cal B}}m^{-\alpha_{\tau}},$ (6.6)

### v2

@@ 4.1 Using FEMs or AFEMs
In order to streamline the discussion using FEM, we make the simplifying but unnecessary assumption that the data tensor product grids $G_{k,r}$ and $\overline{G}_{\overline{k},r}$ are matching, i.e. $k=\overline{k}$. For $n\geq k$, we let $\mathcal{T}_{n}$ be the Kuhn-Tucker subdivision of $\Omega$ associated with the tensor grid $G_{n,r}$ and $V_{n}$ be the spaces of continuous piecewise polynomials of degree $r$ subordinate to $\mathcal{T}_{n}$. We use Lemma 4.1 to derive the following statement about FEMs.
Lemma 4.2.
Let $\widetilde{u}$ be the solution to (1) with right hand side $\widehat{f}:=S^{*}_{k}(f)$ and boundary value $\widehat{g}:=\overline{S}_{k}(g)$. For $n\geq k$, let $\widehat{u}_{n}$ be the Galerkin projection of $\widetilde{u}$ onto $V_{n}$. Then, we have
$\|u-\widehat{u}_{n}\|_{H^{1}(\Omega)}\lesssim R_{\widetilde{m},\overline{m}}^{*}({\cal U})_{H^{1}(\Omega)}+\varepsilon_{n}({\cal U})_{H^{1}(\Omega)},\quad\hbox{where}\quad\varepsilon_{n}({\cal U})_{H^{1}(\Omega)}:=\mathop{\rm dist}({\cal U},V_{n})_{H^{1}(\Omega)}.$ (4.4)
Proof.
It follows from (4.2) that
$\|u-\widehat{u}_{n}\|_{H^{1}(\Omega)}\leq\|u-\widetilde{u}\|_{H^{1}(\Omega)}+\|\widetilde{u}-\widehat{u}_{n}\|_{H^{1}(\Omega)}\lesssim R_{\widetilde{m},\overline{m}}^{*}({\cal U})_{H^{1}(\Omega)}+\|\widetilde{u}-\widehat{u}_{n}\|_{H^{1}(\Omega)}.$ (4.5)
Note that since $n\geq k$, we have that $\widehat{g}\in Tr(V_{n})$. In particular, $\widehat{u}_{n}$ and $\widetilde{u}$ have the same boundary data (and are defined with the same right hand side $\widehat{f}$). Consequently, $\widehat{u}_{n}$ is a quasi-best approximation to $\widetilde{u}$ in $V_{n}$,
$\|\widetilde{u}-\widehat{u}_{n}\|_{H^{1}(\Omega)}\lesssim\inf_{v\in V_{n}}\|\widetilde{u}-v\|_{H^{1}(\Omega)}.$ (4.6)
Invoking (4.2), we find that for every $v\in V_{n}$,
$\|\widetilde{u}-v\|_{H^{1}(\Omega)}\leq\|\widetilde{u}-u\|_{H^{1}(\Omega)}+\|u-v\|_{H^{1}(\Omega)}\lesssim R_{\widetilde{m},\overline{m}}^{*}({\cal U})_{H^{1}(\Omega)}+\|u-v\|_{H^{1}(\Omega)},$
and hence
$\inf_{v\in V_{n}}\|\widetilde{u}-v\|_{H^{1}(\Omega)}\lesssim R_{\widetilde{m},\overline{m}}^{*}({\cal U})_{H^{1}(\Omega)}+\inf_{v\in V_{n}}\|u-v\|_{H^{1}(\Omega)}\leq R_{\widetilde{m},\overline{m}}^{*}({\cal U})_{H^{1}(\Omega)}+\varepsilon_{n}({\cal U})_{H^{1}(\Omega)}.$
This, together with (4.6), gives
$\|\widetilde{u}-\widehat{u}_{n}\|_{H^{1}(\Omega)}\lesssim R_{\widetilde{m},\overline{m}}^{*}({\cal U})_{H^{1}(\Omega)}+\varepsilon_{n}({\cal U})_{H^{1}(\Omega)}.$ (4.7)
Combining (4.5) and (4.7) yields the desired estimate (4.4). ∎
Let us now discuss the statement of Lemma 4.2. We have determined the optimal recovery rates $R_{\widetilde{m},\overline{m}}^{*}({\cal U})_{H^{1}(\Omega)}$, see (3.26) for the class ${\cal U}$. The term $\varepsilon_{n}$ will tend to zero as $n\to\infty$ and we would obtain the optimal recovery error on the right side of (4.4) if $n$ is chosen suitably large. The actual decay rate of $\varepsilon_{n}$ depends on the regularity of the model class ${\cal U}$ in $H^{1}(\Omega)$. Therefore, the issue becomes what do our model class assumptions $f\in{\cal F}$, $g\in{\cal G}$, say about the regularity of $u$ in the $H^{1}(\Omega)$ norm. If we use linear spaces such as standard finite element spaces, then we would want to know the regularity of $u$ in the scale of $H^{t}(\Omega)$ spaces, $t>1$. There are several theorems that obtain such regularity results (see e.g. [12, 42, 55] and the references therein) and thereby obtain concrete bounds on the decay rate $\varepsilon_{n}$. We do not elaborate on this further but refer the interested reader to the papers cited above.
If we use nonlinear numerical methods such as AFEMs for solving (1) with right side $\widehat{f}$, and boundary value $\widehat{g}$, then the decay rate for $\varepsilon_{n}({\cal U})$ can be improved (see [6] for a summary of such results).
@@ 4.2 Reduced models
If one is faced with solving several PDEs with the same data sites $({\cal X},{\cal Z})$ but with different data $({\bf f},{\bf g})$, one can build a numerical algorithm as follows. In the case when ${\cal X}=G_{k,r}$, $|G_{k,r}|=\widetilde{m}=[r2^{k}]^{d}$, see (2.15) and ${\cal Z}=\overline{G}_{\bar{k},r}$, $|\overline{G}_{\bar{k},r}|=\overline{m}=2d[r2^{\bar{k}}]^{d-1}$, see (3.15), we know that the simplicial interpolation operator $S_{k}^{*}$ defined in (2.18) provides a near optimal recovery for ${\cal F}$. Namely,
@@
Similarly, the interpolant $\overline{S}_{\bar{k}}(g)$ to $g\in{\cal G}$ used for optimal recovery, see (3.17), can be written as
$\overline{S}_{\bar{k}}(g)=\sum_{i=1}^{\overline{m}}g({\bf z}_{i})\psi_{i},\quad\overline{m}=2d[r2^{\bar{k}}]^{d-1}.$ (4.10)
@@
is a near optimal recovery of the solution $u$ to (1) with right side $f$ and boundary values $g$. Note that $\widetilde{u}$ is the solution to (1) with right side $S_{k}^{*}(f)$ and boundary value $\overline{S}_{\bar{k}}(g)$.
@@
$\widehat{u}:=\sum_{i=1}^{\widetilde{m}}f({\bf x}_{i})\widehat{\phi}_{i}+\sum_{i=1}^{\overline{m}}g({\bf z}_{i})\widehat{\psi}_{i},\quad\hbox{with}\quad\widetilde{m}\asymp\overline{m}\asymp m/2,$ (4.12)
@@ 5 Numerical methods based on optimization
Another method of finding a near optimal approximation to $u$ from the given data is through optimization, such as the one used in PINNs. The starting point of this approach is the fact that the unique solution $u\in H^{1}(\Omega)$ to (1) is the unique minimizer of the theoretical loss ${\cal L}_{T}(u)$, see (1.6). However, the minimization in (1.6) is over all $v\in H^{1}(\Omega)$. To make this approach more amenable to numerical implementation, we take this minimization over a set $\Sigma=\Sigma_{n}$, where $\Sigma$ is either a linear space of dimension $n$ or a nonlinear manifold of order $n$ (i.e., depending on $n$ parameters), and study any function
@@
$\|u-S_{\Sigma}\|_{H^{1}(\Omega)}\lesssim{\cal L}_{T}(S_{\Sigma})\lesssim{\cal L}_{T}(S)\lesssim\|u-S\|_{H^{1}(\Omega)}.$ (5.2)
Since (5.2) holds for all $S\in\Sigma$, we have that
@@ 6 Discretizing norms
The optimization procedure of the previous section is not a numerical algorithm since it does not incorporate numerical methods for estimating the norms appearing in the loss ${\cal L}_{T}$, namely $\|\cdot\|_{H^{-1}(\Omega)}$ and $\|\cdot\|_{H^{1/2}(\partial\Omega)}$. In this section, we address this issue by introducing discrete norms which involve only the values of functions at the data sites. Replacing the norms in ${\cal L}_{T}$ by these discrete norms leads to a loss ${\cal L}^{*}$ which can be numerically computed. The equivalence between the norms in ${\cal L}_{T}$ and their discrete counterparts holds (modulo corresponding optimal recovery rates) for functions whose Laplacians are uniformly bounded in $\cal B$ and whose traces are uniformly bounded in $\overline{\cal B}$.
As it will be seen in §7, the new loss ${\cal L}^{*}$ is based on a discretization of the $\|\cdot\|_{L_{\gamma}(\Omega)}$ (for a special value of $\gamma$) and $\|\cdot\|_{H^{1/2}(\partial\Omega)}$ norms. These discretizations have to satisfy the following two requirements:
  • use the given information $({\bf f},{\bf g})$ at the grids (data sites) $G_{k,r}\subset\overline{\Omega}$ and $\overline{G}_{\bar{k},r}=G_{\bar{k},r}\cap\partial\Omega$, respectively.
  • the errors incurred for functions in ${\cal F}$ and ${\cal G}$ be comparable with the optimal recovery rates for these classes.
We remark that the discretization of integral norms has been studied in many other papers. For example, in [66] the discretization of the variational form of the Poisson equation for functions in the Barron space (or variation space corresponding to shallow ReLUk networks) is considered. This analysis is done for both random data sites and a fixed grid. There are also papers that prove existence of data sites ${\cal X}$ for which the continuous and discrete $L_{\tau}$ norms are equivalent for elements of certain linear spaces, see [16, 15]. To the best of our knowledge, there are no results in the literature which treat the discretization of the $H^{1/2}$-norm that are comparable to what we do in §6.2.
@@ 6.1 A discrete $L_{\tau}$ norm
In this section, we study the discretization of the $L_{\tau}$ norm for any $1\leq\tau\leq\infty$. Let ${\cal F}=U(B_{q}^{s}(L_{p}(\Omega)))$, $s>d/p,$ be a model class assumption on $f$. We consider the uniform grid $G_{k,r}\subset\overline{\Omega}$, see (2.15), consisting of $\widetilde{m}$ points, $\widetilde{m}=|G_{k,r}|=[r2^{k}]^{d}$, $k\geq 0$ and $r>\max(s,1)$. For any continuous function $f$ and any $1\leq\tau<\infty$, we define
$\|f\|^{*}_{L_{\tau}}:=\left[\frac{1}{\widetilde{m}}\sum_{j=1}^{\widetilde{m}}|f({\rm\bf x}_{j})|^{\tau}\right]^{1/\tau},\quad\hbox{where}\quad{\bf x}_{j}\in G_{k,r}.$ (6.1)
@@
$\|f\|_{L_{\tau}(\Omega)}\lesssim\|f\|^{*}_{L_{\tau}}+\|f\|_{{\cal B}}{\widetilde{m}}^{-\alpha_{\tau}},\quad\hbox{and}\quad\|f\|^{*}_{L_{\tau}}\lesssim\|f\|_{L_{\tau}(\Omega)}+\|f\|_{{\cal B}}{\widetilde{m}}^{-\alpha_{\tau}},$
with $\alpha_{\tau}:=\frac{s}{d}-\left(\frac{1}{p}-\frac{1}{\tau}\right)_{+}$. Recall that ${\widetilde{m}}^{-\alpha_{\tau}}$ is the uniform rate of optimal recovery of the model class ${\cal F}$ in $L_{\tau}(\Omega)$. The constant in $\lesssim$ depends only on $r$ and $d$.
@@
We prove the lemma for $1\leq\tau<\infty$. A simple modification of the proof handles the case $\tau=\infty$. We use the simplicial interpolation operator $S^{*}_{k}$ given in (2.18). Let ${\cal T}_{k}:={\cal T}_{k}(\Omega)=\bigcup T$ be the set of simplices that make up the simplicial decomposition of the cubes in ${\cal D}_{k}(\overline{\Omega})$. Then, from (2.18) we have
@@
where $L_{T}(f)$ is the polynomial in ${\cal P}^{d}_{r}$ that interpolates the data $\{f({\bf x}_{i}):\,{\bf x}_{i}\in\overline{T}\}$. By equivalence of norms on ${\cal P}^{d}_{r}$ (the linear space of algebraic polynomials of $d$ variables and order $r$ (total degree $r-1$)), rescaling arguments, and the fact that we have $d!$ simplices in the Kuhn-Tucker decomposition of a cube,
$\|L_{T}(f)\|_{L_{\tau}(T)}\asymp\Bigl{[}\frac{1}{\widetilde{m}}\sum_{{\bf x}_{i}\in{\overline{T}}}|f({\bf x}_{i})|^{\tau}\Bigr{]}^{1/\tau},$ (6.3)
@@
$\|S_{k}^{*}(f)\|_{L_{\tau}(\Omega)}\asymp\Bigl{[}\frac{1}{\widetilde{m}}\sum_{i=1}^{\widetilde{m}}|f({\bf x}_{i})|^{\tau}\Bigr{]}^{1/\tau}=\|f\|^{*}_{L_{\tau}(\Omega)}.$ (6.4)
For $f\in{\cal B}=B_{q}^{s}(L_{p}(\Omega))$, the piecewise polynomial interpolant $S_{k}^{*}(f)$ provides a near optimal recovery of $f$ from the data ${\bf f}$, that is, see Theorem 2.1,
$\|f-S_{k}^{*}(f)\|_{L_{\tau}(\Omega)}\leq\begin{cases}C|f|_{B_{p}^{s}(\Omega)}2^{-k(s-d/p+d/\tau)}=C\|f\|_{{\cal B}}{\widetilde{m}}^{-\alpha_{\tau}},\quad\quad\quad\quad\quad\quad\tau\geq p,\\ \\ \|f-S_{k}^{*}(f)\|_{L_{p}(\Omega)}\leq C|f|_{B_{p}^{s}(\Omega)}2^{-ks}=C\|f\|_{{\cal B}}{\widetilde{m}}^{-\alpha_{\tau}},\quad\tau<p,\end{cases}$ (6.5)
where we have used that $\widetilde{m}=[r2^{k}]^{d}$. It follows then from (6.4) that
$\left|\|f\|_{L_{\tau}(\Omega)}-\|S_{k}^{*}(f)\|_{L_{\tau}(\Omega)}\right|\leq C\|f\|_{{\cal B}}{\widetilde{m}}^{-\alpha_{\tau}},$ (6.6)